# A quantum Monte Carlo algorithm for arbitrary high-spin Hamiltonians

### v1

@@ Abstract
We present a universal parameter-free quantum Monte Carlo algorithm for simulating arbitrary high-spin (spin greater than $1/2$) Hamiltonians. This approach extends a previously developed method by the authors for spin-$1/2$ Hamiltonians [Phys. Rev. Research 6, 013281 (2024)]. To demonstrate its applicability and versatility, we apply our method to the spin-$1$ and spin-$3/2$ quantum Heisenberg models on the square lattice. Additionally, we detail how the approach naturally extends to general Hamiltonians involving mixtures of particle species, including bosons and fermions. We have made our program code freely accessible on GitHub.
@@ I Introduction
Despite the importance of high-spin models in condensed matter physics, numerical simulation methods for studying these systems are generally scarce (for exceptions addressing specific models, see Refs. [5, 6, 7]). Specifically, quantum Monte Carlo (QMC) techniques, which are essentially the only viable method for studying the finite-temperature equilibrium properties of large-scale quantum many-body systems without approximations, have not yet been fully adapted to the simulation of high-spin quantum systems.
@@
This work aims to remedy this situation by providing a universal framework for studying arbitrarily complex high-spin quantum models of any spin value and interactions of essentially any geometry, dimension, locality, and connectivity. Furthermore, a natural generalization of the proposed framework to address QMC simulations of mixed-species models including spin-$1/2$ particles, bosons, and fermions is also discussed in detail.
@@
The present technique builds on a recent study by the authors [8] in which a universal QMC algorithm designed to simulate arbitrary spin-$1/2$ Hamiltonians was devised. There, it was shown that casting the to-be-simulated Hamiltonian in Permutation Matrix Representation (PMR) form [9] allows one to write the partition function of any spin-$1/2$ system as a sum of efficiently calculable terms each of which is associated with a closed walk on the Hamiltonian graph and where Monte Carlo updates to faithfully sample these walks may be automatically generated in a systematic way [9, 8].
@@
In this paper, we generalize the aforementioned technique to the case of high-spin (spin greater than $1/2$) Hamiltonians and introduce a similar-in-spirit universal Monte Carlo algorithm designed to reliably simulate arbitrary high-spin systems. To that aim, we devise a protocol for generating the necessary set of QMC updates, based on PMR decomposition, that ensure an ergodic Markov chain Monte Carlo sampling of the partition function of essentially any conceivable input system. We illustrate that while for spin-$1/2$ systems achieving the same goal required finding the nullspace of sets of binary (modulo 2) vectors representing the permutation matrices of the Hamiltonian, for spin-$s$ particles the task is generalized to finding a similar set of modulo $(2s+1)$ vectors.
@@
The paper is organized as follows. In Sec. II, we provide a brief overview of the permutation matrix representation quantum Monte Carlo (PMR-QMC) and analyze high-spin Hamiltonians in this context. In Sec. III, we discuss the QMC algorithm, describing in detail the method we have devised to generate all the necessary QMC updates and proving that these moves ensure both ergodicity and detailed balance. In Sec. IV we showcase the power of our technique by presenting simulation results for two models, namely, the spin-$1$ and spin-$3/2$ quantum Heisenberg models on the square lattice, and in Sec. V we discuss in detail how the approach taken can be extended to include other particle species as well as mixtures thereof. We conclude in Sec. VI with an additional discussion and future directions of research.
@@ II.1 Overview of PMR-QMC
where $D_{j}$ are diagonal matrices and $\{P_{j}\}_{j=1}^{M}$ are permutation matrices that are a subset of a special Abelian group, $G$  [^1] Every $P\in G$ is a permutation matrix with no nonzero diagonal elements, except for the identity matrix $P_{0}={\mathbb{1}}$. For every $|i\rangle,|j\rangle$ pair, there exists a unique $P\in G$ such that $P|i\rangle=|j\rangle$. See [15] for more details.. For spin-1/2 systems, $G$ consists of all Pauli-X strings [8].
@@
[…]
@@
The values of $\operatorname{Re}[A_{\cal C}w_{\cal C}]/W_{\cal C}^{(1)}$ and $\left|A_{\cal C}\right|\cos(\arg(A_{\cal C}w_{\cal C}))$, therefore, represent the instantaneous quantity associated with the configuration ${\cal C}={(z,{S_{{\bf{i}}_{q}}})}$ that will be collected during the simulation when using the weights $W_{\cal C}^{(1)}$ and $W_{\cal C}^{(2)}$, respectively. For further details about calculating a wide range of observables throughout the simulation, see Refs. [9, 8, 15].
@@ II.2 PMR decomposition of single spin operators
Before discussing the PMR formulation of general high-spin Hamiltonians, let us first briefly review high-spin operators. The matrix elements of the $(2s+1)\times(2s+1)$ spin matrices $X$, $Y$, and $Z$ for $s\in\{1/2,1,3/2,2,5/2,\ldots\}$ are given by
@@
where $j,k\in\{1,2,\dots,2s+1\}$. In the spirit of PMR, these matrices can be written in the form
@@
respectively, where $P$ is the $(2s+1)\times(2s+1)$ permutation matrix given by
$P=\begin{pmatrix}0&1&0&\ldots&0\\ 0&0&1&\ldots&0\\ \vdots&\vdots&\ddots&\ddots&0\\ 0&0&\ldots&0&1\\ 1&0&\ldots&0&0\\ \end{pmatrix},$ (13)
which obeys $P^{2s+1}=\mathbb{1}$, and $D^{+}$, $D^{-}$, $D^{(z)}$ are diagonal matrices with the following diagonal entries:
@@
For reasons that will become clear shortly, we also introduce the matrices $D^{(k)}$ and $D^{(z,k)}$ such that
@@
where $D^{+}_{0}$ and $D^{(z)}_{0}$ are defined as $D^{+}_{0}=D^{+}_{2s+1}=0$ and $D^{(z)}_{0}=D^{(z)}_{2s+1}=-s$. We note that $D^{+}=D^{(0)}$, $D^{-}=D^{(-1)}$, and $D^{(z)}=D^{(z,0)}$. The matrices $D^{(k)}$ and $D^{(z,k)}$ obey
@@ II.3 High-spin Hamiltonians
where $c_{i}$ are real-valued coefficients and $\prod_{k=1}^{m_{i}}s^{(i)}_{j_{i,k}}$ are spin operator strings. Here, $s^{(i)}_{j_{i,k}}$ represents a spin matrix $s\in\{X,Y,Z\}$ in the $i$-th string acting on the $j_{i,k}$-th particle, where $j_{i,k}\in\{1,2,\dots,n\}$. The operator $s^{(i)}_{j_{i,k}}$ is a tensor product of a spin matrix and $n-1$ identity matrices such that it has the same matrix dimension as the Hamiltonian. Each spin operator string may contain any number of any of the three spin matrices for each of the particles in any order.
@@ III.2 Fundamental cycles
Denoting by $p_{i}$ the integer-string $[n_{i,1}n_{i,2}\cdots n_{i,n}]$, one can easily verify that the product of two permutation operators $P_{i}$ and $P_{k}$ would likewise correspond to modular addition of the components of $p_{i}$ and $p_{k}$ modulo $12s+$ 1, i.e.
@@
For a sequence of operators evaluating to the identity, $S_{{\bf{i}}_{q}}=P_{i_{q}}\ldots P_{i_{1}}$, we have $\sum_{j=1}^{q}p_{i_{j}}\equiv{\bf 0}\,(\text{mod }2s+1)$, where ${\bf 0}$ is an integer-string consisting of only zeros. We note that $S_{{\bf{i}}_{q}}$ is a permutation of the multiset of operators $\{P_{i_{1}},\dots,P_{i_{q}}\}$, which can be represented as an integer string $[a_{1}\,a_{2}\,\dots\,a_{M}]$, where $a_{k}$ denotes the number of occurrences of the operator $P_{k}$ among $P_{i_{1}},\dots,P_{i_{q}}$.
The question of how one can generate all possible sequences of operators (which evaluate to the identity) is therefore reduced to the question of how one can generate all integer-strings $[a_{1}\,a_{2}\,\dots\,a_{M}]$ that obey the following system of linear equations over mod-$(2s+1)$ addition
@@
Equation (30) can be solved by finding the null space basis over addition modulo $12s+$ 1 for the matrix ${\bf n}^{T}$ whose columns are $p_{1}^{T},\ldots,p_{M}^{T}$. For details on accomplishing this task regardless of the decomposition of $12s+$ 1 into prime factors, the reader is referred to Appendix B.
@@
It follows from the properties of null space that any cycle can be obtained by (i) insertion and removal of fundamental cycles, (ii) insertion and removal of trivial cycles consisting of $12s+$ 1 identical permutation operators, and (iii) swapping the order of two adjacent permutation operators.

### v2

@@ Abstract
We present a universal quantum Monte Carlo algorithm for simulating arbitrary high-spin (spin greater than $1/2$) Hamiltonians, based on the recently developed permutation matrix representation (PMR) framework. Our approach extends a previously developed PMR-QMC method for spin-$1/2$ Hamiltonians [Phys. Rev. Research 6, 013281 (2024)]. Because it does not rely on a local bond decomposition, the method applies equally well to models with arbitrary connectivities, long-range and multi-spin interactions, and its closed-walk formulation allows a natural analysis of sign-problem conditions in terms of cycle weights. To demonstrate its applicability and versatility, we apply our method to spin-$1$ and spin-$3/2$ quantum Heisenberg models on the square lattice, as well as to randomly generated high-spin Hamiltonians. Additionally, we show how the approach naturally extends to general Hamiltonians involving mixtures of particle species, including bosons and fermions. We have made our program code freely accessible on GitHub.
@@ I Introduction
In the literature, a variety of approaches have been proposed to tackle numerous large-scale high-spin models, ranging from a functional renormalization group approach to study Heisenberg models with unrestricted spin length [5] to a high-order coupled cluster method applied to Kagome lattice antiferromagnets with arbitrary spin restricted to translationally invariant lattices [6]. Also notable are studies investigating spin-liquid behavior in a spin-one Kitaev model under magnetic fields [7] which are closely tied to the unique solvability of the Kitaev model.
The main workhorse in the study of high-spin quantum systems is quantum Monte Carlo (QMC) techniques. The stochastic series expansion (SSE) with operator-loop updates [8] has been successful for simulating Heisenberg models. Directed loop algorithms [9] extended the SSE approach providing improved ergodicity and efficiency, albeit their implementation for high spins remain Hamiltonians remain model-specific. Additionally, there have been numerous QMC proposals to address the sign-problem for frustrated spin systems. For example, in Ref. [10] a QMC scheme for frustrated Heisenberg antiferromagnets has been developed to mitigate the sign-problem in frustrated systems. Other studies (see, e.g., Ref. [11]) explored mixed-spin quantum magnets where spin magnitudes vary across sites, and the finite-temperature behavior of square-lattice spin-one Heisenberg antiferromagnets [12].
As the above literature survey suggests, high-spin models are of significant relevance in condensed matter physics; however, existing numerical techniques for simulating high-spin quantum many-body systems remain constrained in both generality and applicability. In particular, QMC algorithms often rely on update schemes that are intricately tailored to the specific symmetries and properties of the model under study, rendering them non-transferable across different systems and limiting their broader utility.
@@
This work aims to remedy this situation by providing in contrast a universal framework for studying arbitrarily complex high-spin quantum models of any spin value and interactions of essentially any geometry, dimension, locality, and connectivity. Furthermore, a natural generalization of the proposed framework to address QMC simulations of mixed-species models including spin-$1/2$ particles, bosons, and fermions is also discussed in detail.
@@
The present technique builds on a recent study by the authors [13] in which a universal QMC algorithm designed to simulate arbitrary spin-$1/2$ Hamiltonians was devised. There, it was shown that casting the to-be-simulated Hamiltonian in Permutation Matrix Representation (PMR) form [14] allows one to write the partition function of any spin-$1/2$ system as a sum of efficiently calculable terms each of which is associated with a closed walk on the computational state graph of the Hamiltonian  [^1] The computational state graph of a Hamiltonian is the graph $G_{H}=(V,E)$ whose nodes are the computational basis states, and the edges (connectivity and weights) are given by the off-diagonal terms of the Hamiltonian $H$. and where Monte Carlo updates to faithfully sample these walks may be automatically generated in a systematic way [14, 13].
The PMR-QMC framework has already been benchmarked in several spin-$1/2$ settings, including transverse-field Ising and XXZ models, the toric code, and random MAX2SAT-type Hamiltonians [16, 14, 13]. In these studies PMR-QMC was found to equilibrate reliably, and in some nonlocal problems, such as random MAX2SAT in a transverse field, it outperformed optimized SSE implementations by orders of magnitude in wall-clock time while yielding results that agree within statistical error.
@@
In this paper, we generalize the aforementioned technique to the case of high-spin (spin greater than $1/2$) Hamiltonians and introduce a similar-in-spirit universal Monte Carlo algorithm designed to reliably simulate arbitrary high-spin systems. To that aim, we devise a protocol for generating the necessary set of QMC updates, based on the PMR decomposition, that ensure an ergodic Markov chain Monte Carlo sampling of the partition function of essentially any conceivable input system. We illustrate that while for spin-$1/2$ systems achieving the same goal required finding the nullspace of sets of binary (modulo 2) vectors representing the permutation matrices of the Hamiltonian, for spin-$s$ particles the task is generalized to finding a similar set of modulo $(2s+1)$ vectors.
@@
The paper is organized as follows. In Sec. II, we provide a brief overview of the permutation matrix representation quantum Monte Carlo (PMR-QMC) and analyze high-spin Hamiltonians in this context. In Sec. III, we discuss the QMC algorithm, describing in detail the method we have devised to generate all the necessary QMC updates and demonstrating how these moves ensure both ergodicity and detailed balance. There, we also discuss the emergence of the sign problem in our scheme. In Sec. IV we showcase the power of our technique by presenting simulation results for two models, namely, the spin-$1$ and spin-$3/2$ quantum Heisenberg models on the square lattice, as well as for randomly generated spin-$1$ and spin-$3/2$ Hamiltonians. In Sec. V we examine in detail how the approach taken can be extended to include other particle species as well as mixtures thereof. We conclude in Sec. VI with an additional discussion and future directions of research.
@@ II.1 Overview of PMR-QMC
where $D_{j}$ are diagonal matrices and $\{P_{j}\}_{j=1}^{M}$ are permutation matrices. As shown in [17], these permutation matrices can be chosen as a subset of a special Abelian group $G$  [^2] Every $P\in G$ is a permutation matrix with no nonzero diagonal elements, except for the identity matrix $P_{0}={\mathbb{1}}$. For every $|i\rangle,|j\rangle$ pair, there exists a unique $P\in G$ such that $P|i\rangle=|j\rangle$. See [17] for more details.. For spin-1/2 systems, $G$ consists of all Pauli-X strings [13]. For $S>\frac{1}{2}$ systems, we will describe a decomposition of spin operators ($SU(2)$ generators), in section II.2, using permutation matrices that are by construction Abelian.
@@
is efficiently computable. Here, ${\bf{i}}_{q}=(i_{1},\ldots,i_{q})$ is a multi-index, where each index $i_{j}$ runs from $1$ to $M$, and $S_{{\bf{i}}_{q}}\equiv P_{i_{q}}\ldots P_{i_{1}}$ denotes a product of $q$ permutations, each from $\{P_{j}\}_{j=1}^{M}$. The summation over $\cal C$ is a double sum over all basis states $|z\rangle$ and all possible products $S_{{\bf{i}}_{q}}$ that evaluate to the identity, for $q$ from 0 to $\infty$. Next, we denote $|z_{0}\rangle\equiv|z\rangle$ and $|z_{k}\rangle\equiv P_{\mathbf{i}_{k}}\ldots P_{\mathbf{i}_{1}}|z\rangle$ for $k=1,2,\dots,q$. This notation allows us to define the ‘diagonal-energies’ as $E_{z_{k}}\equiv\left<z_{k}|H|z_{k}\right>=\left<z_{k}|D_{0}|z_{k}\right$> and the off-diagonal ‘hopping strength,’ $D_{(z,S_{{\bf{i}}_{q}})}\equiv\prod_{k=1}^{q}d^{({\bf{i}}_{k})}_{z_{k}}$, where $d^{({\bf{i}}_{k})}_{z_{k}}=\left<z_{k}|D_{i_{k}}|z_{k}\right$>. Finally, $e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}$ denotes the divided difference [19, 14] of $f(x)=e^{-\beta x}$ with respect to the inputs $\{E_{z_{0}},\ldots,E_{z_{q}}\}$, which can be efficiently computed in practice in $O(q)$ time [20].
At the formal level, as expressed in Eqs. (2) and (3), the PMR-QMC framework can be viewed as a reorganization of the conventional stochastic series expansion (SSE) representation. Like Handscomb’s original diagrammatic expansions [21, 22] and Sandvik’s SSE algorithm [8], PMR-QMC starts from the Taylor expansion of $e^{-\beta H}$. However, instead of sampling explicit operator strings built from local bond terms, PMR-QMC groups the Taylor terms into closed walks of basis states and expresses their weights as divided differences of the exponential. In this representation PMR-QMC is no longer a high-temperature expansion, and a single PMR-QMC configuration aggregates an entire family of SSE operator strings that differ only in the ordering and multiplicity of diagonal insertions.
The validity of the PMR construction does not depend on a particular choice of computational basis: given any orthonormal basis of the Hilbert space, one can construct the corresponding permutation matrices and run PMR-QMC in that basis. Although the explicit forms of the matrices change from one basis to another, the structure of the algorithm and its Monte Carlo updates remains the same. This basis flexibility is a useful practical feature, as it allows one to perform local basis rotations which, in some cases, can mitigate or even remove a sign problem by mapping the Hamiltonian into a more favourable form (cf. Sec. III.4).
@@
The values of $\operatorname{Re}[A_{\cal C}w_{\cal C}]/W_{\cal C}^{(1)}$ and $\left|A_{\cal C}\right|\cos(\arg(A_{\cal C}w_{\cal C}))$, therefore, represent the instantaneous quantity associated with the configuration ${\cal C}={(z,{S_{{\bf{i}}_{q}}})}$ that will be collected during the simulation when using the weights $W_{\cal C}^{(1)}$ and $W_{\cal C}^{(2)}$, respectively. For further details about calculating a wide range of observables throughout the simulation, see Refs. [14, 13, 17, 24]. In particular, this formalism allows one to construct estimators for essentially arbitrary static operators as well as for imaginary-time correlation functions and their integrated susceptibilities. The explicit examples of such constructions given in Refs. [17, 24] are fully applicable to the high-spin case.
@@ II.2 PMR decomposition of single spin operators
Before discussing the PMR formulation of general high-spin Hamiltonians, let us first briefly review high-spin operators. The matrix elements of the $(2s+1)\times(2s+1)$ spin operators $X$, $Y$, and $Z$ for $s\in\{1/2,1,3/2,2,5/2,\ldots\}$ are given by
@@
where $j,k\in\{1,2,\dots,2s+1\}$.
In order to devise a PMR decomposition for these spin operators, let us define the following $(2s+1)\times(2s+1)$ permutation matrix
$P=\begin{pmatrix}0&1&0&\ldots&0\\ 0&0&1&\ldots&0\\ \vdots&\vdots&\ddots&\ddots&0\\ 0&0&\ldots&0&1\\ 1&0&\ldots&0&0\\ \end{pmatrix}$. (10)
We must note that $P^{2s+1}=\mathbb{1}$ and that $P$ has no fixed points on the computational basis states (eigenstates of the spin-$Z$ operator). Using this permutation matrix, we can write the spin operators as follows
@@
where $^{+}, DD^{-}$, $D^{(z)}$ are diagonal matrices with the following diagonal entries:
@@
For reasons that will become clear in the following sections, we also introduce the matrices $D^{(k)}$ and $D^{(z,k)}$ such that
@@
where $D^{+}_{0}$ and $D^{(z)}_{0}$ are defined as $D^{+}_{0}=D^{+}_{2s+1}=0$ and $D^{(z)}_{0}=D^{(z)}_{2s+1}=-s$. We note that $^{+}=D^{(0)}, D=D$$^{-}=D^{(-1)}$, and $D^{(z)}=D^{(z,0)}$. The matrices $D^{(k)}$ and $D^{(z,k)}$ obey
@@
We highlight the features of this PMR decomposition through the following example. Consider the following two-spin operator $Y\otimes X$ and let the spin $s>1/2$. The PMR decomposition for this will be
$Y\otimes X$ $=(-iD^{+}P+iD^{-}P^{-1})\otimes(D^{+}P+D^{-}P^{-1})$
$=-i(D^{+}P\otimes D^{+}P+D^{+}P\otimes D^{-}P^{-1})$
$\,\,\,\,\,+i(D^{-}P^{-1}\otimes D^{+}P+D^{-}P^{-1}\otimes D^{-}P^{-1})$
$=-i\Big((D^{+}\otimes D^{+})P_{1}+(D^{+}\otimes D^{-})P_{2}$
$\quad-(D^{-}\otimes D^{+})P_{3}-(D^{-}\otimes D^{-})P_{4}\Big)\,$, (23)
where $P_{1}=P\otimes P$, $P_{2}=P\otimes P^{-1}$, $P_{3}=P^{-1}\otimes P$, and $P_{4}=P^{-1}\otimes P^{-1}$. We emphasize that for spin-$1/2$, $P=P^{-1}=X$, so there will only be a single unique permutation operator as $P_{1}=P_{2}=P_{3}=P_{4}$. However, this is not that case for higher spin ($s>1/2$) systems. Additionally, one can verify that these permutation operators commute, i.e. $P_{i},P_{j}]=0$ for $i,j\in\{1,2,3,4\}$, highlighting the fact that tensor product of powers of single-spin permutation operators $P$ are Abelian and can be used for PMR decomposition for any spin Hamiltonian.
@@ II.3 High-spin Hamiltonians
where $c_{i}$ are real-valued coefficients and $\prod_{k=1}^{m_{i}}s^{(i)}_{j_{i,k}}$ are spin operator strings. Here, $s^{(i)}_{j_{i,k}}$ represents a spin matrix $s\in\{X,Y,Z\}$ in the $i$-th string acting on the $j_{i,k}$-th particle, where $j_{i,k}\in\{1,2,\dots,n\}$. The operator $s^{(i)}_{j_{i,k}}$ is a tensor product of a spin matrix and n1n-1 identity matrices such that it has the same matrix dimension as the Hamiltonian. Each spin operator string may contain any number of any of the three spin matrices for each of the particles in any order.
@@ III.2 Fundamental cycles
Denoting by $p_{i}$ the integer-string $n_{i,1}n_{i,2}\cdots n_{i,n}]$, one can easily verify that the product of two permutation operators $P_{i}$ and $P_{k}$ would likewise correspond to modular addition of the components of $p_{i}$ and $p_{k}$ modulo $2s+1$, i.e.
@@
For a sequence of operators evaluating to the identity, $S_{{\bf{i}}_{q}}=P_{i_{q}}\ldots P_{i_{1}}$, we have $\sum_{j=1}^{q}p_{i_{j}}\equiv{\bf 0}\,(\text{mod }2s+1)$, where $\bf 0$} is an integer-string consisting of only zeros. We note that $S_{{\bf{i}}_{q}}$ is a permutation of the multiset of operators $\{P_{i_{1}},\dots,P_{i_{q}}\}$, which can be represented as an integer string $a_{1}\,a_{2}\,\dots\,a_{M}]$, where $a_{k}$ denotes the number of occurrences of the operator $P_{k}$ among $P_{i_{1}},\dots,P_{i_{q}}$.
The question of how one can generate all possible sequences of operators (which evaluate to the identity) is therefore reduced to the question of how one can generate all integer-strings $a_{1}\,a_{2}\,\dots\,a_{M}]$ that obey the following system of linear equations over mod-$(2s+1)$ addition
@@
Equation (31) can be solved by finding the null space basis over addition modulo $2s+1$ for the matrix ${\bf n}^{T}$ whose columns are $p_{1}^{T},\ldots,p_{M}^{T}$. For details on accomplishing this task regardless of the decomposition of $2s+1$ into prime factors, we utilize the fact that a system of modular equivalences Eq. (31) can be rewritten as a system of Diophantine equations
where $k_{1},k_{2},\dots,k_{n}$ are additional unknown integers, and $s_{1}=\dots=s_{n}=s$. The Diophantine system Eq. (32) can be solved by employing Hermite normal form of the corresponding matrix. For details, the reader is referred to Refs. [27, 28].
@@
It follows from the properties of null space that any cycle can be obtained by (i) insertion and removal of fundamental cycles, (ii) insertion and removal of trivial cycles consisting of $2s+1$ identical permutation operators, and (iii) swapping the order of two adjacent permutation operators.